OpenStudy (nesha97):

For the given polynomials, complete the following tasks: • List the possible rational roots. • Express the polynomial in factored form. x^3-x^2+4x-4 3x^4-5x^3-5x^2+5x+2

OpenStudy (nesha97):

can someone please show how to do this for the first one i think the possible rots are 1,2,4 for the second one i think its 1/2, 1, 2 not sure how to do the rest....please help

OpenStudy (nesha97):

for the first one i mean negative and positive for 1,2,4

OpenStudy (irishboy123):

there's a grouping technique which is listed here for cubics in practise it is worth just guessing and using long division for the second trickier one \(3x^4-5x^3-5x^2+5x+2\) start with x = 1 \(3-5-5+5+2 = 0\) Bingo!! then with x = -1 \(3+5-5-5+2 = 0\) Bingo!! \( \dfrac {3x^4-5x^3-5x^2+5x+2} {x^2 -1 } = 3x^2 - 5x -2\) and apply the quadratic formula unless, unlike me, you can see further roots \(\dfrac{5 \pm \sqrt{ (-5)^2 - 4(3)(-2) }}{2(3)}\) \(= \dfrac{5 \pm \sqrt{ 25 + 24 }}{6}\) gives you \(-\frac{1}{3}, 2\) is that what you meant?!?!

OpenStudy (nesha97):

so -1/3, 2 are my rational roots?

OpenStudy (osprey):

@Nesha97 @IrishBoy123 since they are fractions, then they seem to represent rational numbers (pi is irrational 'cos you can't express it as a fraction), I'd guess that yes is the answer to your q.

OpenStudy (osprey):

the fact that 1/3 is 0.33333333333333333333333 until the end of space-time-ink-paper is slightly embarrasing .... leave it as 1/3 ?

OpenStudy (nesha97):

how do i express the polynomial in factored form?

OpenStudy (irishboy123):

for 4 roots, \(r_1, r_2, r_3, r_4\) its \(f(x) = (x-r_1)(x-r_2)(x-r_3)(x-r_4)\) And yes, whole numbers are rational so it's the whole shebang

OpenStudy (irishboy123):

\(f(x) = (x-1)(x \color{red}{+}1)(x \color{red}{+} \frac{1}{3})(x-2)\)

OpenStudy (nesha97):

okay thank you

OpenStudy (irishboy123):

mp